OpenStudy (anonymous):

write 5*16+7*16 in the simplest form

OpenStudy (anonymous):

\[\LARGE a*b+c*a=a(b+c)\]

OpenStudy (anonymous):

the multiplications are done first = 80 + 112

OpenStudy (anonymous):

192...

OpenStudy (anonymous):

(5x16) + (7x16) (80) + (112) = 192 :)

OpenStudy (anonymous):

what kind of simple do they mean?

OpenStudy (anonymous):

good question ! LOL

OpenStudy (inkyvoyd):

This is a perfect example, @apoorvk . This is an example of the one medal thing I was talking about.

OpenStudy (inkyvoyd):

@apoorvk ?

OpenStudy (inkyvoyd):

Look at who got medals. Tell me one medal system is fair.

OpenStudy (apoorvk):

yes, post the link to this in the feedback page, and do it under a new question if you like, this would really help to draw attention to the matter.

OpenStudy (inkyvoyd):

Can you do it? I'm a noooob!

OpenStudy (apoorvk):

or just post a screenshot, and after that give medals to the right guys here. what say?

OpenStudy (anonymous):

What's the problem with the medals? I got one, so it's a good system.

OpenStudy (inkyvoyd):

Yes. Yes it is *sarcasm end*

OpenStudy (anonymous):

I guess you know better than most who is the one that deserves the medals.

OpenStudy (apoorvk):

@scarydoor I 'm afraid this isn't the way how it works. This is not an answering site, it's a helping website. You help and explain the concepts, not just post an answer for people to copy down. :)

OpenStudy (anonymous):

@apoorvk I'm not sure what you're afraid of. Poster above me posted an equation. Question was "write in simplest way possible" I figured the next most simple thing to a+b, which had already been given, would be the result of a+b, rather than repeating again the derivation of a+b which had already been covered.

OpenStudy (inkyvoyd):

This is not wikianswers, or yahoo answers. THis is a study group, we teach, we learn. We don't cheat, nor do we just give out answers.

OpenStudy (inkyvoyd):

@scarydoor

OpenStudy (anonymous):

im still not sure how you got the answer..my son seems to think 12/16 and 6/16 are the answers now i have been out of school for years and math was never my best subject so can someone explain please

OpenStudy (anonymous):

the answer is 192 lol

OpenStudy (anonymous):

@inkyvoyd I'm sorry that I told the questioner what 112 + 80 is equal to. Shame on me!

OpenStudy (apoorvk):

yes, i think so too @scarydoor because the asker just said that he still isn't clear about this, although he got the answer a long time back.

OpenStudy (anonymous):

Then what are you doing not explaining it to him/her? There are different ways of helping someone with maths. I was running on the assumption that the asker knows what is meant by simplification, and that me explaining the various small technicalities in the addition of two numbers would only bore the asker. If the asker does not understand, then they can ask more questions. I suggest that if you are so concerned about whether the questioner properly grasps the solution, you would be adding to an explanation directed toward that person, rather than criticizing my technique which is to demonstrate that further addition of numbers is possible, and that this does lead to a simpler solution. This is called teaching through demonstration and is a well documented and acknowledged form of teaching.

OpenStudy (anonymous):

@terrance basically the question just wants you to rearrange the expression so that it looks more simple. 5*16 + 7*16 contains all the information you need to get one single number. There are rules in maths about which operations are performed before others. read this: http://en.wikipedia.org/wiki/Order_of_operations in maths multiplication is performed before addition. So, first compute 5*16 and then compute 7*16. Then you can add the result of that together and you'll get 192. Alternatively you could rearrange 5*16 + 7*16 into (5+7)*16. Think of 5*16 as a grid which has length 5 and length 16. 5*16 is what you get if you count each square. Similarly for 7*16. Now 5*16 + 7*16 is the result of counting all squares in two grids. One has sides 5 and 16, the other has sides 7 and 16. If you join the two together you can get one grid that has sides (5+7) and the other side is 16. So now you have (5+7)*16. That is one way to justify why 5*16 + 7*16 = (5+7)*16.

OpenStudy (anonymous):

As for why your son got the answers 12/16 and 6/16 I am not sure. Have you typed the question correctly above? Or maybe he has mixed up the meaning up the symbols, or performed operations in the incorrect order. Above, do you mean that * is to stand for multiplication?

OpenStudy (inkyvoyd):

Try using the draw option to give us what the problem looks like, and we'll try to help out some more.

OpenStudy (inkyvoyd):

And @scarydoor , giving the answer directly is never encouraged on openstudy in any case, showing work then getting an answer is acceptable, but giving the answer is never acceptable for a number of reasons mentioned in the CoC.

OpenStudy (inkyvoyd):

If you really still disagree with my statement, just tag a moderator; I'm sure they will be glad to explain.

OpenStudy (anonymous):

@inkyvoyd I did not know that I am your servant. If you want to draw the diagram, then be my guest. However I believe my description is sufficient.

OpenStudy (inkyvoyd):

And, I was talking to @terrance , I do apologize for the misunderstanding I may have caused you, @scarydoor .

OpenStudy (anonymous):

@inkyvoyd if I see someone attempting to simplify a problem and leaving it in the form 112+80, then I won't hesitate to complete the solution as 192. If you want to say that is a problem, then go ahead with it. But I think that's crazy.

OpenStudy (apoorvk):

@scarydoor now your explanation deserved a medal. and you got it from me. wish you understood this earlier.

OpenStudy (anonymous):

I'm glad that I measure up to your standards. *laugh*